OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally 1,250 cubic meters and is then resized using a scale factor of one fifth on each side length?

OpenStudy (anonymous):

ok, so you have an inital volume that is of 1250, which is l * h * w. so here's what we'll do : 1250 = l * w* h your new volume is this : (l/5) * (w/5) * (h/5) = X combine all the coefficients : (1/5*1/5*1/5) * (l*w*h) = X now we already found out that l * w* h =1250, so we can replace it in this equation : (1/5*1/5*1/5) * 1250 = X (1/125) * 1250 = X 10 = X Bottom line, simply cube your length ratio factor ((1/5)^3) to get the volume ratio

OpenStudy (anonymous):

thank youu so much ! (:

OpenStudy (anonymous):

yw